OpenStudy (anonymous):

How fast would a 15 kg box accelerate down a 25 degree snowy slope if the coefficient of kinetic friction were 0.15?

OpenStudy (anonymous):



OpenStudy (anonymous):

sum your forces in the direction PARALLEL with the incline: = friction force - gravity of box = u*mg*cos(25) - mg*sin(25) = 0.15*15*9.81*cos(25) - 15*9.81*sin(25) = -42.18 now, we know the box is moving, so that's why we never said "0=" when writting the equation above., so, we use a "-ma " negative becuase its moving down, ma becuase f=ma, and if we have the -42.18 as a unit of force, the other side of the equation must also be a unit of force., so we have: -ma = -42.18 -15*a = -42.18 a = 42.18/15 boo ya